the roots of the eqation (x-1)^3+8=0 are

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear manan

(x-1)^3+8=0

by inspection we can easly find out that x=-1 is the root of equation

now f(x) = (x-1)^3+8

      f'(x) = 3(x-1)2

  f'(x) 0 for all x

so f(x) is a increasing function

so given function can cut x axix only 1 point ,equation has only one root = -1

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin